OpenStudy (anonymous):

How to simplify? x+2 ---- x-2

OpenStudy (anonymous):

Do you mean \[\frac{ x+2 }{ x-2 }\] ?

OpenStudy (anonymous):

Yes, that's correct

OpenStudy (mathstudent55):

There is nothing to simplify.

OpenStudy (anonymous):

Hmm, it says evaluate as well. Does that change anything?

OpenStudy (mathstudent55):

No it isn't -1.

OpenStudy (anonymous):

2 doesn't work either.

OpenStudy (mathstudent55):

If you were asked to evaluate it, were you given an x value to evaluate it at?

OpenStudy (anonymous):

No, it just says: Evaluate and simplify. Assume that no denominator is zero.

OpenStudy (mathstudent55):

\(\dfrac{ x+2 }{ x-2 }\) Let x = 5: \(\dfrac{ 5+2 }{ 5-2 } = \dfrac{7}{3} \) Here's a single counterexample that shows this is not -1. This expression will have a different value for each value of x, and it is not defined for x = 2 since x = 2 will cause the denominator to be zero.

OpenStudy (anonymous):

Right yes, I was thinking of multiplying it by the converse for some reason ><

OpenStudy (mathstudent55):

@nincompoop I'm sure you meant \(x \ne 2\), not \(n \ne 0\).

OpenStudy (dumbcow):

you could evaluate the graph of the function i guess. vertical asymptote at x=2 horizontal asymptote at y=1 http://www.wolframalpha.com/input/?i=plot+%28x%2B2%29%2F%28x-2%29

OpenStudy (nincompoop):

yes x not equal to 2

OpenStudy (nincompoop):

I shouldn't even try to help right now ... laughing out loud my brain is fried

OpenStudy (anonymous):

When you set x - 2 =0 We get x = 2 Therefore, since we cannot divide by 0, x cannot equal 2 It is all real numbers where \[x \neq 2\] But this equation cannot be simplified further. Which is what it is asking. Right?

OpenStudy (mathstudent55):

Correct. As I stated in my first response, there is nothing you can do to simplify this.

OpenStudy (anonymous):

Okay, I looked at this question and had no idea what to do. I guess there isn't anything haha. Thanks you guys.

OpenStudy (anonymous):

Rewrite the problem as the answer. There is no simplification to be made, so just write it again as the solution.